I need to write an equation of the line containing the given points and parallel to the given line. Express answer in y=mx+b forn
(-9,6); 2x=3y+8 The equation of the line is y=

Thank you

Found 3 solutions by mananth, haileytucki, Alan3354:
Answer by mananth(16949)

I need to write an equation of the line containing the given points and parallel to the given line. Express answer in y=mx+b forn
(-9,6); 2x=3y+8 The equation of the line is y=

2x=3y+8
re write in the form y = mx +b
3y=2x-8
/3
y =(2x/3)-(8/3)
slope = m = 2/3
...
The slope of the required line will be the same since the line is parallel.
y= mx + b
(-9,6) the point & slope = 2/3
Find b by plugging thses values
6=(2/3)*-9+b
6=-6+b
b=12
..
slope = 2/3, b =12
The equation will be y = (2x/3)+12
...

m.ananth@hotmail.ca

Answer by haileytucki(390)

(-9,6)_2x=3y+8
Since y is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
3y+8=2x
Since 8 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 8 from both sides.
3y=-8+2x
Move all terms not containing y to the right-hand side of the equation.
3y=2x-8
Divide each term in the equation by 3.
(3y)/(3)=(2x)/(3)-(8)/(3)
Simplify the left-hand side of the equation by canceling the common factors.
y=(2x)/(3)-(8)/(3)
To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept.
y=mx+b
Using the y=mx+b formula, m=(2)/(3).
m=(2)/(3)
The negative reciprocal of I is 0.
mperp=-(3)/(2)
Find the equation of the perpendicular line using the point-slope formula.
(-9,6)_m=-(3)/(2)
Find the value of b using the formula for the equation of a line.
y=mx+b
Substitute the value of m into the equation.
y=(-(3)/(2))*x+b
Substitute the value of x into the equation.
y=(-(3)/(2))*(-9)+b
Substitute the value of y into the equation.
(6)=(-(3)/(2))*(-9)+b
Since b is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
(-(3)/(2))*(-9)+b=(6)
Multiply (-(3)/(2)) by (-9) to get (-(3)/(2))(-9).
(-(3)/(2))(-9)+b=(6)
Remove the parentheses around the expression 6.
(-(3)/(2))(-9)+b=6
Multiply -(3)/(2) by -9 to get (27)/(2).
((27)/(2))+b=6
Reorder the polynomial (27)/(2)+b alphabetically from left to right, starting with the highest order term.
b+(27)/(2)=6
Find the value of b.
b=-(15)/(2)
Now that the values of m(slope) and b(y-intercept) are known, substitute them into y=mx+b to find the equation of the line.
y=-(3x)/(2)-(15)/(2)

Answer by Alan3354(69443)

A line and a point example.
Write in standard form the eqation of a line that satisfies the given conditions. Perpendicular to 9x+3y=36, through (1,2)
-----------------
Find the slope of the line. Do that by putting the equation in slope-intercept form, y = mx + b. That means solve for y.
9x+3y = 36
3y= - 9x + 36
y = -3x + 12
The slope, m = -3
------------------
The slope of lines parallel is the same.
The slope of lines perpendicular is the negative inverse, m = +1/3
----------------
Use y = mx + b and the point (1,2) to find b.
2 = (1/3)*1 + b
b = 5/3
The equation is y = (1/3)x + 5/3 (slope-intercept form)
x - 3y = -5 (standard form)
------------------------
For further assistance, or to check your work, email me via the thank you note, or at Moral Loophole@aol.com